Performance Tuning Scienti�c Codes for Data�ow Execution

Computation Structures Group Memo 381 October 19, 1996

Andrew Shaw

Arvind

R. Paul Johnson

Appeared in Parallel Architectures and Compiler Techniques ’96

This report describes research done at the Laboratory for Computer Science and the Department of Earth, Atmospheric, and Planetary Sciences of the Massachusetts Institute of Technology. Funding for the Laboratory for Computer Science is provided in part by the Advanced Research Projects Agency of the Department of Defense under the Of�ce of Naval Research contract N00014-92-J-1310.

Copyright 1996 IEEE. Published in Parallel Architectures and Compilation Techniques, October 20-23, 1996, Boston, MA, USA. Personal use of this material is permitted. However, permission to reprint/republish this material for advertising or promotional purposes or for creating new collective works for resale or redistribution to servers or lists, or to reuse any copyrighted component of this work in other works, must be obtained from the IEEE. Contact: Manager, Copyrights and Permissions / IEEE Service Center / 445 Hoes Lane / P.O. Box 1331 / Piscataway, NJ 08855-1331, USA. Telephone: + Intl. 908-562-3966.

1